For equation of y^2dy= x(xdy-ydx)e^(x/y)

1. Find the function M(x,y) and its degree
2. Find the function N(x,y) and its degree
3. Is the given Differential Equation a homogenous? If yes, what is the degree?
1
Expert's answer
2021-09-16T06:57:01-0400

Solution;

Rewrite the equation as follows;

"y^2dy=(x^2dy-xydx)e^{\frac xy}"

"y^2dy=x^2e^{\frac xy}dy-xye^{\frac xy}dx"

Combine as follows;

"xye^{\frac xy}dx=(x^2e^{\frac xy}-y^2)dy" ...(a)

1.)

From the above equation (a),

"M(x,y)=xye^{\frac xy}dx"

The degree of M(x,y) is 1.

2.)

From the above equation (a);

"N(x,y)=x^2e^{\frac xy}-y^2"

The degree of N(x,y) is 2.

3.)

Because of "e^{\frac xy}" ,take "\frac{dx}{dy}" instead of "\frac{dy}{dx}"

From equation (a);

We obtain;

"\frac{dx}{dy}=\frac{x^2e^{\frac xy}-y^2}{xye^{\frac xy}}"

Now put ,F(x,y)="\frac{dx}{dy}" and find "F(\lambda x,\lambda y)"

Hence we have ;

"F(\lambda x,\lambda y)=\frac {(\lambda x)^2e^{\frac{\lambda x}{\lambda y}}-(\lambda y)^2}{(\lambda x)(\lambda y)e^{\frac{\lambda x}{\lambda y}}}"

"F(\lambda x,\lambda y)=\frac{(\lambda )^2[x^2e^{\frac xy}-y^2]}{\lambda^2xye^{\frac xy}}" =F(x,y)

Clearly;

"F(\lambda x,\lambda y)=F(x,y)=\lambda ^0F(x,y)"

Therefore ,the given equation is a homogeneous function of degree zero(0).